# revealing an in-context neural scaling law

### v1

@@ Abstract
Pretrained large language models (LLMs) are surprisingly effective at performing zero-shot tasks, including time-series forecasting. However, understanding the mechanisms behind such capabilities remains highly challenging due to the complexity of the models. In this paper, we study LLMs’ ability to extrapolate the behavior of dynamical systems whose evolution is governed by principles of physical interest. Our results show that LLaMA 2, a language model trained primarily on texts, achieves accurate predictions of dynamical system time series without fine-tuning or prompt engineering. Moreover, the accuracy of the learned physical rules increases with the length of the input context window, revealing an in-context version of neural scaling law. Along the way, we present a flexible and efficient algorithm for extracting probability density functions of multi-digit numbers directly from LLMs.
@@ 1 Introduction
Large language models (LLMs) are billion-parameter neural networks that have been trained on a vast corpus of text to perform next-word predictions (Radford et al., 2019; Brown et al., 2020; Kaplan et al., 2020; OpenAI, 2023; Touvron et al., 2023). Since the introduction of the transformer architecture (Vaswani et al., 2017), LLMs have become remarkably proficient in a wide range of natural language processing (NLP) tasks, including question answering (Roberts et al., 2020; Drori et al., 2022) and translation (Brown et al., 2020). In the process of training larger models, unexpected emergent properties of LLMs have been discovered (Wei et al., 2022a), such as program execution (Nye et al., 2021) and multi-step reasoning (Cobbe et al., 2021; Wei et al., 2022b; Suzgun et al., 2022).
While powerful, LLMs remain scientifically mysterious. How they compress vast amounts of information and implement complex algorithms within their architecture is not readily discoverable. To attempt to shine some light on these “black box” models, recent studies notably used representation probes to decipher some of the mechanisms encoded in the layers of trained neural networks (Alain & Bengio, 2016; Akyürek et al., 2022; Gurnee & Tegmark, 2023; Marks & Tegmark, 2023; Nanda et al., 2023; Park et al., 2023). By showing, experimentally, that features of prompt representations in a transformer’s latent space can be mapped onto human-comprehensible concepts (e.g., statement veracity, geographical localization, etc.), these findings suggest the possibility of an internal “world model”, which provides LLMs with some sort of perception of reality, far more sophisticated than that of a mere “stochastic parrot.”
In this work, we continue these explorations of intrinsic world models by proposing a new perspective and empirical approach. Specifically, we are interested in the (apparent) unreasonable effectiveness of LLMs in understanding the physical world. We then leverage and extend recent observations that LLMs are capable of time series extrapolation without specific prompting or fine-tuning (Gruver et al., 2023; Jin et al., 2023a). Instead of forecasting empirical time series, we study the ability of LLMs to extrapolate dynamical systems time series while maintaining fidelity with underlying, governing principles.
Figure 1: Continuing a non-linear pendulum angular trajectory using LLaMA-13b. The prediction accuracy increases with the input context length. The temperature parameter is set to $10^{-7}$ to eliminate randomness in auto-regressive continuation.
As an example, Figure 1 illustrates an LLM’s ability to learn in-context and continue the deterministic trajectory of a non-linear pendulum, whose amplitude and frequency both drift over time (Section A.4). The prediction accuracy improves with the length of the input time series, which we refer to as context length.
We focus on systems that are either deterministic or stochastic but have well-defined transition rules between time steps. We compare the transition rules inferred by the LLM with the ground truth underlying the input data by using specific loss functions. We find that loss functions decrease rapidly with the number of states observed in-context, revealing a neural scaling law for in-context learning, as illustrated in Figure 2. This observed behavior spans a wide range of dynamical systems, from physically-grounded systems like the Lorenz system to entirely synthetic ones, such as Markov chains with arbitrary transition probabilities. Our results suggest that LLMs might possess not only an internal model of the world (Gurnee & Tegmark, 2023), but moreover an in-context learning ability which seems to allow them to glean physical insights from input data.
@@
Organization of the paper. We begin in Section 2 by providing background information on LLMs for time series forecasting. We then describe our methods for tokenizing time series and extracting probability density functions from LLMs in Section 3. Our experiments and their results on various dynamical systems using the LLaMA model (Touvron et al., 2023) are presented in Section 4. We conclude in Section 5 by discussing the implications of our findings for the understanding of in-context learning abilities and the intrinsic world model of LLMs.
Main contributions. The main contributions of this paper are summarized as follows:
@@
  • discovering a scaling law between the accuracy of the learned transition rules and the context window length.
@@ 2 Background and related work
LLMs are primarily designed, trained, and utilized for NLP, i.e. the ability to generate meaningful language in response to a textual prompt. However, scaled-up modern models have been found to perform surprisingly well for tasks beyond NLP (Wei et al., 2022a).
Among these emergent capabilities, a specific class called in-context learning, consists of learning from and extrapolating sequences included in the prompt, even when they are not language (Brown et al., 2020). One useful example of in-context learning is zero-shot time series forecasting (Gruver et al., 2023). Given a numerical time series of sufficient length, LLMs can extrapolate the prompt sequence and generate numerical values that are consistent with the input without any fine-tuning, and sometimes more accurately than specialized algorithms for time series extrapolation (Wu et al., 2021; Zhou et al., 2022).
@@
Unlike these prior studies, our paper does not focus on forecasting real-world time series, such as weather data (Zhou et al., 2021) or electricity demand (Godahewa et al., 2021), where the underlying model generating the sequence is unavailable or undefined. Instead, we aim to extract the learned transition rules from the probability vector generated by the LLM and compare them against the ground truth rules (chaotic, stochastic, discrete, continuous, etc.) governing the input time series.
@@ 3 Methodology
    Measure the discrepancy between the ground truth $P_{ij}$ and learned $\tilde{P}_{ij}$ using appropriate loss functions.
@@ 3.1 Prompt generation
Markov processes. Most of our testing data may be modeled as discrete-time Markov chains, where the PDF of the next state at time $t+1$ depends solely on the previous state $x_{t}$ at time $t$:
$P(X_{t+1}|X_{1}=x_{1},\dots,X_{t}=x_{t})=P(X_{t+1}|X_{t}=x_{t}).$
This models either discrete iterative systems or continuous dynamical systems after time-discretization using a first-order explicit time-stepping scheme. Each time series is governed by a set of transition rules $P(X_{t+1}|X_{t},\Delta t)$, which, in general, depends on both the previous state $x_{t}$ and the time step $\Delta t$. However, we assume a uniform time-discretization of the system (i.e., $\Delta t$ is constant), and consider time-independent transition rules: $P(X_{t+1}|X_{t})$. Hence, we are interested in how LLMs learn transition rules in-context, as opposed to forecasting unseen transition rules. It is worth noting that the Markovian nature does not imply that the transformer only relies on $x_{t}$ to predict $X_{t+1}$. Instead, an efficient model might combine all previous states $\{x_{0},\ldots,x_{t}\}$ to hypothesize a reasonable transition rule $P(X_{t+1}|X_{t})$, which can then be used, in conjunction with the last observed state $x_{t}$ to formalize a probabilistic prediction of $X_{t+1}$.
Time series tokenization. An input time series typically consists of ($\sim 10^{3}$) time steps, each represented as a real number. We first rescale each number and represent it using $n$ digits (typically, $n=3$). Each time series is rescaled to the interval $[1.50,8.50]$ so that the number of digits never changes throughout the series. In particular, there is an intentional gap between the upper limit $8.5$ and $10$, so that if an LLM outputs $1$ as the first digit, it is to be unambiguously interpreted as a number $1.xx$, as opposed to $1x.x$. Similarly, the lower limit of $1.5$ further eliminates ambiguity by circumventing leading zeros.
We then follow the scheme introduced in (Gruver et al., 2023; Jin et al., 2023a) to serialize the time series as strings and tokenize them. Since our numerical experiments are performed with LLaMA models (Touvron et al., 2023) (though our conclusions apply to other models, see Appendix A), which contain ten distinct tokens for the digits $0,\ldots,9$, we pass the string-digits without spacing but with a comma separating time steps. In the top right of Figure 4, we show an example of a time series serialized as an input string.
@@ 3.2 Extraction of transition rules
Discrete state space. When the Markov process is discrete and has a finite state space, each state can be represented by a single token. While one could theoretically select any subset within the set of 32k tokens in the LLaMA codebook, we employ tokens corresponding to the ten number strings: $0,\ldots,9$. Hence, these strings are represented with ten independent tokens, while letter tokens are often combined to form sub-components of words. Moreover, we observe in Section 4 that the most sophisticated LLaMA model (LLaMA-70b) that we evaluate can only learn up to 9 discrete states. Therefore, we do not attempt to go beyond 9 distinct states by extending to non-number tokens.
@@
Figure 3 illustrates our framework for learning discrete Markov chains with LLMs. First, we randomly draw an $n\times n$ matrix $(P_{ij})$ from the space of transition matrices. We then sample the Markov chain, tokenize the series and prompt the LLM with it. The length of the series is chosen such that the tokenized representation does not exceed the length of the LLM’s context window. We extract the LLM’s prediction for the next state by performing a softmax operation on the output logits corresponding to the $n$ allowed states and discarding all other logits. This is equivalent to modifying the non-state logits using an additive bias of negative infinity, which is consistent with the sampling method described in (Gruver et al., 2023; Jin et al., 2023a).
@@
Continuous-time state space. Stochastic processes, such as the Brownian motion (Einstein, 1905; Perrin, 1909), might have continuous-time state space $\{X_{t}\}_{t\geq 0}$, indexed by a real variable $t\geq 0$. For these processes, we represent the value of each state as a multi-digit number and separate each state using the comma symbol “,”, which also has its own token in the LLaMA codebook, and is never lumped along with neighboring numbers. As observed in (Jin et al., 2023a), an LLM prediction of multi-digit values can be naturally interpreted as a hierarchical softmax distribution (Mnih & Hinton, 2008; Challu et al., 2023).
Figure 4: Extracting transition rules of systems with continuous state space. The learned transition rules are more refined for bins near the mode, which dominate the Bhattacharyya distance. The height of $\tilde{P}(x_{t+1}|x_{t})$ is shown in log scale.
@@
Specifically, let $u$ denote a multi-digit string representing the value of a state at a given time-step, then the LLM’s softmax prediction for the $i^{th}$ digit, $u_{i}$, provides a histogram of ten bins of width $0.1^{i}$. Subsequently, the prediction of the $(i+1)^{\mathrm{th}}$ digit goes down one level into the hierarchical tree by refining one of the bins into ten finer bins of width $0.1^{i+1}$, and so on until the last digit is processed (see Figure 4).
@@
Hierarchy-PDF algorithm. While a single pass through the LLM yields a discretized PDF represented by bins of various widths, we can refine the PDF by querying each coarse bin. For example, to obtain a maximal resolution PDF of a 3-digit value, we need to query all $10^{2}$ combinations of the first two digits and furnish a PDF of uniform bin width, $0.1^{3}$. Suppose a time series consists of S values (steps), each represented as $n$ digits, then furnishing maximal precision PDFs for the entire sequence requires $10^{n-1}S$ forward passes of the LLM. This seemingly daunting complexity is significantly reduced because most of the $10^{n-1}S$ inputs differ only in the last tokens, and thus one can recursively cache the key and value matrices associated with the shared tokens. The complexity is further reduced by refining only the high-probability bins near the mode, as shown in Figure 4. Algorithm 1 outlines the Hierarchy-PDF algorithm used to recursively refine the PDF associated with a multi-digit value in a time series. More algorithmic details are available in Appendix A as well as in the associated GitHub repository.
@@
        Change the last digit to launch 9 recursive branches
        RecursiveRefiner($D_{c}$, $D_{t}$)
@@
        refine multi_PDF with new logits
@@
           Append digits to launch 10 recursive branches
           RecursiveRefiner($D_{c}+1$, $D_{t}$)
@@
3.3 Loss functions
@@
where the mean $\mu_{\tilde{P}}=\sum_{x}\tilde{p}(x)x\Delta x$ is extracted from the hierarchical PDF. Note that unlike the Bhattacharyya distance, which references the model prediction $\tilde{p}$ only at $x_{\mathrm{true}}$, the SDM takes into account the entire support $x\in\mathcal{X}$. Our numerical experiments suggest that SDM is more stable and better captures the in-context learning dynamics of deterministic systems (see Section 4.1).
4 Experiments and Results
This section reports empirical in-context learning results on several dynamical systems modeling physics phenomena, such as a Brownian motion, the logistic map with chaotic behavior, and the Lorenz system. All experiments are repeated with ten trajectories initiated by different random seeds for stochastic processes, or initial conditions for chaotic systems. In each case, we extract the transition rules $P(X_{t}|X_{t-1})$ learned by the model, and then calculate the in-context loss curves, using the Hierarchy-PDF algorithm described in Section 3.
Choice of the model. All numerical experiments reported in this section are performed using the LLaMA-13b model (Touvron et al., 2023), which is open source, unlike GPT-4, and does not have the tokenization issue associated with GPT-3 (Gruver et al., 2023). While we observe that larger language models such as LLaMA-70b may achieve lower in-context loss on some dynamical systems (Section A.5), they do not display qualitative differences.
4.1 Deterministic Systems
4.1.1 Logistic map
$x_{t+1}=f(x_{t})=rx_{t}(1-x_{t}),\quad x_{0}\in(0,1),$ (3)
which may also be written using conditional distributions to reflect the deterministic nature of the system as
$X_{t+1}|\{X_{t}=x_{t}\}\sim\delta_{f(x_{t})},$
where $\delta$ denotes the Dirac delta distribution. This conditional distribution is the ground truth transition function displayed in red in Figure 5. The parameter $r\in[1,4)$ controls the behavior of the system and is set to $r=3.9$. At this value, the dynamics are naturally confined within the interval $(0,1)$, and the system has no stable fixed points. Due to the chaotic nature of the system, two initial nearby trajectories diverge exponentially in time. This property allows us to sample multiple uncorrelated trajectories by using different initial conditions $x_{0}$, sampled uniformly in $(0,1)$.
4.1.2 Lorenz system
$\dot{x}(t)$ $=\sigma(y-x),\quad\dot{y}(t)=x(\rho-z)-y,$
$\dot{z}(t)$ $=xy-\beta z,$
4.2 Stochastic Systems
4.2.1 Brownian motion

### v2

@@ Abstract
Pretrained large language models (LLMs) are surprisingly effective at performing zero-shot tasks, including time-series forecasting. However, understanding the mechanisms behind such capabilities remains highly challenging due to the complexity of the models. We study LLMs’ ability to extrapolate the behavior of dynamical systems whose evolution is governed by principles of physical interest. Our results show that LLaMA 2, a language model trained primarily on texts, achieves accurate predictions of dynamical system time series without fine-tuning or prompt engineering. Moreover, the accuracy of the learned physical rules increases with the length of the input context window, revealing an in-context version of neural scaling law. Along the way, we present a flexible and efficient algorithm for extracting probability density functions of multi-digit numbers directly from LLMs.
@@ 1 Introduction
Since the introduction of the transformer architecture (Vaswani et al., 2017), Large language models (LLMs) have shown a variety of unexpected emergent properties, such as program execution (Nye et al., 2021) and multi-step reasoning (Cobbe et al., 2021; Wei et al., 2022; Suzgun et al., 2022).
Despite the empirical success of LLMs, how they compress vast amounts of information and implement complex algorithms within their architecture is not readily discoverable. To this end, recent studies used representation probes to decipher how concepts and functions are encoded in the layers of trained neural networks (Akyürek et al., 2022; Gurnee & Tegmark, 2023; Marks & Tegmark, 2023; Park et al., 2023; Hendel et al., 2023).
@@
Our work explores LLMs’ ability to model the world by proposing a new perspective and empirical approach. Inspired by the recent observations that LLMs are capable of in-context time series extrapolation without specific prompting or fine-tuning (Gruver et al., 2023; Jin et al., 2023a), we aim to quantify LLMs’ ability to extrapolate stochastic dynamical systems. We find that, as the number of observed time steps increases, an LLM’s statistical prediction consistently converges to the ground truth transition rules underlying the system; leading to an empirical scaling law, as observed in Figure 1.
Our main contributions are as follows:
@@
  • discovering a scaling law between the accuracy in the learned transition rules (compared to ground truth) and the context window length.
@@ 2 Background and related work
In-context learning refers to LLM’s emergent ability to learn from examples included in the prompt (Brown et al., 2020). One example of in-context learning is zero-shot time series forecasting (Gruver et al., 2023).
@@
Unlike these prior studies, our work does not focus on forecasting real-world time series, such as weather data or electricity demand, where the underlying model generating the sequence is unavailable or undefined. Instead, we aim to extract the learned transition rules from the probability vector generated by the LLM and compare them against the ground truth rules (chaotic, stochastic, discrete, continuous, etc.) governing the input time series.
@@ 3 Methodology
    Measure the discrepancy, between the ground truth $P_{ij}$ and learned $\tilde{P}_{ij}$, using Bhattacharyya distance.  [^1] Other loss functions may be appropriate depending on whether the dynamical system is stochastic or deterministic, see Section A.1
@@ 3.1 Prompt generation
Markov processes. Most of our testing data may be modeled as discrete-time Markov chains, where the probability distribution function (PDF) of the next state at time $t+1$ depends solely on the previous state $x_{t}$ at time $t$:
$P(X_{t+1}|x_{1},\dots,x_{t})=P(X_{t+1}|x_{t}).$
This models either discrete iterative systems or continuous dynamical systems after time-discretization.
Time series tokenization. An input time series typically consists of ($\sim 10^{3}$) time steps, each represented as a real number. We first rescale each number and represent it using $n$ digits (typically, $n=3$). Each time series is rescaled to the interval $[1.50,8.50]$ so that the number of digits never changes throughout the series. We then follow the scheme introduced in (Gruver et al., 2023) to serialize the time series as strings and tokenize them.
@@ 3.2 Extraction of transition rules
Discrete state space. When the Markov process is discrete and has a finite state space, each state can be represented by a single token. We employ tokens corresponding to the ten number strings: $0,\ldots,9$. We find that even the most sophisticated LLaMA model (LLaMA-70b) can only learn up to 9 discrete states. Therefore, we do not attempt to go beyond 9 distinct states by extending to non-number tokens (see Section A.3.3 for more details).
@@
Figure 2 illustrates our framework for learning discrete Markov chains with LLMs. First, we randomly sample an $n\times n$ transition matrix $(P_{ij})$. We then generate a Markov chain according to $P_{ij}$, tokenize the time series and pass to an LLM with no additional “prompt engineering”. The length of the series is chosen such that the tokenized representation does not exceed the length of the LLM’s context window. We extract the LLM’s prediction for the next state by performing a softmax operation on the output logits corresponding to the $n$ allowed states and discarding all other logits.
@@
Continuous state space. Many stochastic processes, such as the Brownian motion (Einstein, 1905; Perrin, 1909), are supported on continuous state space. For these processes, we represent the value of each state as a multi-digit number and separate each state using the comma symbol “,”. As observed in (Jin et al., 2023a), an LLM prediction of multi-digit values can be naturally interpreted as a hierarchical softmax distribution (Mnih & Hinton, 2008; Challu et al., 2023).
@@
Specifically, let $u$ denote a multi-digit string representing the value of a state at a given time-step, then the LLM’s softmax prediction for the $i^{th}$ digit, $u_{i}$, provides a histogram of ten bins of width $0.1^{i}$. Subsequently, the prediction of the $(i+1)^{\mathrm{th}}$ digit goes down one level into the hierarchical tree by refining one of the bins into ten finer bins of width $0.1^{i+1}$, and so on until the last digit is processed (see Figure 3). The top right of the figure shows an example of a time series serialized as an input string.
Figure 3: An example of hierarchical transition rules extracted from LLaMa-13b. The PDF bins are color-coded based on resolutions, which in this example are more refined near the mode. The height of $\tilde{P}(x_{t+1}|x_{t})$ is shown in log scale.
@@
Hierarchy-PDF algorithm. While a single pass through the LLM yields a discretized PDF represented by bins of various widths, we can refine the PDF by querying each coarse bin. For example, to furnish a maximal resolution PDF of a 3-digit number, we need to query all $10^{2}$ combinations of the first two digits of that number. Suppose a time series consists of $S$ values (steps), each represented as $n$ digits. Obtaining a maximal resolution PDF for each value of the entire sequence requires $10^{n-1}S$ forward passes of the LLM. This daunting process could be significantly simplified because most of the $10^{n-1}S$ inputs differ only in the last tokens, and thus one can recursively cache the key and value matrices associated with the shared tokens. The computation can be further reduced by refining only the high-probability bins near the mode, which dominate the loss functions, as shown in Figure 3. Algorithm 1 outlines the Hierarchy-PDF algorithm used to recursively refine the PDF associated with a multi-digit value in a time series (more details available in  Appendix A).
@@
        Alter the last digit to launch 9 recursive branches
        RecursiveRefiner(PDF, $D_{c}$, $D_{t}$)
@@
        refine PDF with new logits
@@
           Append the last digit to launch 10 recursive branches
           RecursiveRefiner(PDF, $D_{c}+1$, $D_{t}$)
@@
  Output: Refined PDF
4 Experiments and Results
This section reports empirical in-context learning results on two example systems: discrete Markov chain and stochastic logistic map. We defer discussion of other systems, reported in Figure 1, to Sections A.3 and A.4. The experiments are repeated ten times with trajectories initiated by different random seeds.
Model choice. All numerical experiments reported in this section are performed using the open-source LLaMA-13b model. While we observe that larger language models, such as LLaMA-70b, may achieve lower in-context loss on some dynamical systems (Section A.3.3), they do not display qualitative differences and affect our conclusions.
4.1 Markov chains with discrete states
$P_{ij}=P(X_{t+1}=j|X_{t}=i),\quad 1\leq i,j\leq n.$
Using the testing procedures elaborated in Section 3.2, we generate 10 Markov chains, each from a distinct and randomly generated transition matrix of size $n=4$.
Figure 4: Markov chain in-context loss curves decay rapidly with respect to the input time series length. The average loss is obtained from 10 individual loss curves.
The corresponding loss curves between the LLM predictions and the ground truth are displayed in Figure 4. The LLM formulates remarkably accurate statistical predictions as more time steps are observed in context, even though the transition rules are synthesized completely at random. These conclusions hold for larger transition matrices ($n>4$) and more sophisticated LLMs, such as LLaMA-70b (see Section A.3.3).
4.2 Noisy logistic map
The logistic map, first proposed as a discrete-time model for population growth, is one of the simplest dynamical systems that manifest chaotic behaviors (Strogatz, 2015). It is governed by the following iterative equation:
$x_{t+1}=f(x_{t})=rx_{t}(1-x_{t}),\quad x_{0}\in(0,1),$
where $r\in[1,4)$ is a parameter. The logistic map system becomes stochastic when one introduces small Gaussian perturbations of variance at each step, resulting in modified iterative equation:
$x_{t+1}=f(x_{t}+\epsilon),$
where $\epsilon\stackrel{{\scriptstyle\text{i.i.d}}}{{\sim}}\mathcal{N}(0,\sigma^{2})$. In this case, the ground truth distribution of the next state, $x_{t+1}$, conditioned on the current state $x_{t}$ is Gaussian with mean $f(x_{t})$ and variance $(\sigma f^{\prime}(x_{t}))^{2}$:
The first derivative of $f$ measures how sensitive the local dynamics are to external perturbations. This intuitively explains why the standard deviation of the conditional distribution should be proportional to $f^{\prime}$. We note that the approximation in Equation 1 assumes a small perturbation compared to the second derivative, that is $\sigma^{2}\ll 1/f^{\prime\prime}(x)$.
Figure 5: Stochastic logistic map in-context loss curves.
We again observe a power-law-like decay of the in-context loss function with respect to the length of the observed time series in Figure 5. To achieve low in-context loss, the LLM must learn to predict not only the mean, but also the variance of future steps. This is shown in Figure 6 and discussed further in Section A.9.
5 Discussion and conclusion
Data leakage. The possibility that LLMs’ accurate predictions of the next time step value are merely due to memorization seems extremely unlikely. A sequence of even one thousand numerical values, encoded with three digits, may cover $10^{3000}$ particular instances, which is well beyond the $\sim 10^{12}$ tokens of the training corpus (Touvron et al., 2023).
In-context neural scaling law. Neural scaling laws (Kaplan et al., 2020) are power laws that characterize how the loss of trained neural networks vary with respect to parameters of the model, such as model size, dataset size, and computational resources. To the best of our knowledge, neural scaling laws have so far only been observed in the training procedure, which updates the weights of neural networks using an explicit algorithm, such as stochastic gradient descent and Adam (Kingma & Ba, 2014). The loss curves observed in the different numerical experiments (see Figures 1 and 18) reveal an additional in-context scaling law for LLMs zero-shot learning of dynamical systems. Further analysis of these scaling laws are resented in Section A.6.
Main conclusions. We showed that, with increasing cardinality in tokenization, and given sufficient context, LLMs can accurately extrapolate not just deterministic trajectories (Section A.4), but also chaotic and stochastic sequences governed by specific transition rules. In the latter case, the extrapolation accuracy is measured in a statistical sense. This suggests that large language models can in-context learn dynamical systems’ time series and predict future states in a manner that maintains fidelity with the underlying principles governing the system’s evolution. Moreover, this behavior is observed without the use of any fine-tuning or “prompt engineering”. Our results suggest that a transformer, trained primarily on textual data, can extract governing principles of numerical sequences observed in-context.
Future directions. The in-context neural scaling law hints at a learning algorithm that LLMs implicitly implement during inference, such as gradient descent (Von Oswald et al., 2023). Characterizing such an algorithm is an open question of broad interest  (Shen et al., 2023). Another exciting future direction is to study the generalization of the observed in-context neural scaling laws for other LLMs, such as GPT4 (OpenAI, 2023), and the newly proposed state space models (Gu & Dao, 2023).
Software and Data
Appendix A Appendix
A.1 Loss Functions
@@
where the mean $\mu_{\tilde{P}}=\sum_{x}\tilde{p}(x)x\Delta x$ is extracted from the hierarchical PDF. Note that unlike the Bhattacharyya distance, which references the model prediction $\tilde{p}$ only at $x_{\mathrm{true}}$, the SDM takes into account the entire support $x\in\mathcal{X}$. Our numerical experiments suggest that SDM is more stable and better captures the in-context learning dynamics of deterministic systems (see Section A.4).
A.2 Additional loss functions
Although commonly used as the training loss for a variety of machine learning systems, this loss function may suffer from numerical instabilities as the learned transition function $\tilde{P}$ are often close to zero, as shown in Figures 13 and 8, where the probability density is concentrated in small regions of the support.
$=-\frac{1}{2}\ln(\tilde{p}(x_{\mathrm{true}}))+\text{constant}.$
As a result, the Bhattacharyya distance is reduced to an affine-transformed negative log-likelihood assigned to data by the LLM. Such local sensitivity on $\tilde{p}(x_{\mathrm{true}})$ explains the wild fluctuations seen in the Bhattacharyya loss in Section A.4.
The kurtosis is equal to $3$ for a Gaussian distribution and $9/5$ for bounded uniform distributions. Figure 7 shows the kurtosis of Brownian motion transition rules learned by LLM, which converges to $3$ as the context length increases.
A.3 Additional Experiments: stochastic time series
A.3.1 Brownian motion